Церковний рiк Воскресiння

34. Христос воскрес!

Старовинний кант Галицький розспiв

[figure: 4 4; 4 4; Хрис-; 2; тос; воск-; 3; рес; iз; 4; мер-; 5; твих,; 6; смер-; тi-; ю]

[figure: 7; смерть; здо-; лав; 8; i; тим,; що; 9; у; гро-; бах,; 10; жит-; тя; да-; ру-; 11; вав!]

†Павло Богмат MusiX TEX